OpenStudy (anonymous):

help me 2^0=1 2^-1=1/2

OpenStudy (anonymous):

do you divide them

OpenStudy (anonymous):

cuz i got 0.5

OpenStudy (anonymous):

Is this a true or false question?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

What is the question?

OpenStudy (anonymous):

Simplify\[simplify\left( 2^{0} \right)\times \left( 2^{-1} \right)\]

OpenStudy (unklerhaukus):

\[x^n\times x^m=x^{n+m}\]

OpenStudy (unklerhaukus):

where does the divide but come in?

OpenStudy (anonymous):

@UnkleRhaukus Does that mean it is \[2^-1\]

OpenStudy (unklerhaukus):

yep

OpenStudy (unklerhaukus):

which equals 1/2=0.5

OpenStudy (anonymous):

Yea :D Your full of ideas